# Quantum Thermodynamic Integrability for Canonical and non-Canonical Statistics

### v3

@@ Abstract
between two distributions. In FIG. 2(c-d), the quantities of non-canonical states are calculated to compare with those of canonical states, which are evaluated both numerically and analytically as functions of the relative deviation $\sigma$. Respectively, FIG 2(c) illustrates the difference in the internal energy $\Delta U\equiv U^{(\mathrm{nc})}-U$ (red circles and curve) and the generalized free energy $\Delta F\equiv F^{(\mathrm{nc})}-F$ (blue squares and curve), and FIG 2(d) shows the difference in information entropy (blue squares and curve) and thermodynamic entropy (red circles and curve) for the non-canonical distributions. In the figure, all markers are evaluated numerically, and the curves represent the leading order. Here we have chosen , $\Delta_{1}=1$, $\Delta_{2}=2$, and $\beta=1$.
Correlation from Non-canonical state - For the above non-canonical states defined by Eq. (10), we can check that $S_{\mathrm{in}}(P_{n},P_{m}^{\prime})\neq S_{\mathrm{in}}(P_{m})+S_{\mathrm{in}}(P_{n})$ for $S_{\mathrm{in}}(P_{n})=-\sum_{n}P_{n}\ln P_{n}$. The information entropy for the subsystems just shows no additivity due to the non-canonicality. Such correlation is quantified by the mutual information $I\equiv S_{\mathrm{in}}(P_{m})+S_{\mathrm{in}}(P_{n})-S_{\mathrm{in}}(P_{n},P_{m}^{\prime})$. We plot the mutual information of two correlated three-level systems as a function of relative deviation in FIG. 3.
@@
Figure 3: Mutual information between two three levels systems as a function of the relative deviation of inverse temperature.
@@
Actually, the non-canonicality of a black hole induces the information embedded in the correlation of its radiation, which has been considered a promising candidate for the black hole information paradox. The question arises whether the non-canonical solutions to EIE can describe such non-canonical radiation field from a small Schwarzchild back hole. For a Schwarzchild black hole with temperature $\beta=8\pi M$, the deviation of $\beta$ is estimated as $\Delta\beta^{2}=\left|\beta^{2}/C\right|=8\pi$, where $C=-8\pi M^{2}$ is the heat capacity of the black hole (Hawking, 1975; Wald, 2001; Landau and Lifshitz, 2013; Carlip, 2014). From non-canonical solutions of EIE, the emission rate (see Sec. V of Supplementary Materials (SM_, )) reads as
@@
which is valid in the cases with small deviation. This result is in agreement with the former results obtained from the microscopic theory for the radiation field near the small black hole (Parikh and Wilczek, 2000). The non-canonical state is capable of accounting for the black hole information paradox (Zhang et al., 2009; Cai et al., 2016). It is emphasized that such a non-canonical solution in this letter is regarded as the result of thermodynamic integrability. As black holes become smaller, the back-reaction of radiation grows increasingly stronger, leading to a tighter correlation between successive radiation events. This heightened correlation causes the radiation state to deviate significantly from the traditional canonical thermal equilibrium state.
Conclusion - In conclusion, we have proposed an alternative approach to reconsider the foundations of statistical mechanics according to the quantum thermodynamic integrability (QTI) that we introduced in this letter as a generalization of the path-independence of work and heat. We established a quantum version of Carathéodory’s principle and derived the so-called entropy integrable equations (EIE), which can result in both canonical and non-canonical states. By considering detailed balance, we uniquely constructed the Boltzmann distribution from EIE for QTI as the foundation of statistical physics. In this sense, QTI not only reinforces the principles of statistical physics based on the experiment-driven principles of thermodynamics, but also provides a new perspective on non-canonical thermodynamics based on information correlation. This work opens up new avenues for addressing longstanding problems in statistical thermodynamics, offering a more comprehensive framework for analyzing complex thermodynamic behaviors in small systems.
Acknowledgment - We are grateful to H. Dong, H. T. Quan and Y. H. Ma for their helpful discussions. This work has been supported by the National Natural Science Foundation of China (NSFC) (Grant No. 12088101) and NSAF No.U2330401.

### v4

@@ Abstract
To quantify the difference between the canonical and noncanonical states, we investigate an example of three-level systems with $E_{0}=0$, $E_{1}=\Delta_{1}$ and $E_{2}=\Delta_{2}$. In FIG. 2, the thermodynamic functions of noncanonical states are calculated to compare with those of canonical states, which are evaluated both numerically and analytically as functions of the relative deviation $\sigma$. Respectively, FIG 2(a) illustrates the difference in the internal energy $\Delta U\equiv U^{(\mathrm{nc})}-U$ (red circles and curve) and the generalized free energy $\Delta F\equiv F^{(\mathrm{nc})}-F$ (blue squares and curve), and FIG 2(b) shows the difference in information entropy (blue squares and curve) and thermodynamic entropy (red circles and curve) for the noncanonical distributions. In the figure, all markers are evaluated numerically, and the curves represent the leading order shown in Sec. IV in Supplementary Material (SM_, ). Here we have chosen $\Delta_{1}=1$, $\Delta_{2}=2$, and $\beta=1$.
Systems in the noncanonical states usually show observable effects, especially for those with phase transitions. For canonical states, thermodynamic functions of a phase transition system exhibit discontinuity at the critical point, and such discontinuity is smoothed for noncanonical states. We take the two dimensional Ising model as an example. In canonical states, the system exhibits a phase transition at the critical temperature $\beta_{c}=\ln(1+\sqrt{2})/2J$, where $J$ is the amplitude of the exchange interaction between spins. Near the critical point, the macroscopic magnetization $M$ is zero for $t<0$ and increases sharply as $M\propto t^{1/8}$ for $t>0$, and the heat capacity $C(\beta)$ diverges as $\ln|t|$ (Onsager, 1944; Yang, 1952). Here $t\equiv(\beta-\beta_{c})/\beta$ is the relative distance to the critical point. Such critical behaviors for canonical states are illustrated in FIG. 3 with black dotted curves.
For the noncanonical states, the thermodynamic functions near $\beta_{c}$ are dependent on $t$ and $\sigma$. We illustrate in FIG. 3 the noncanonical magnetization and heat capacity, demonstrating that the critical behaviors are smoothed. The data for $\sigma=0.02$ and $\sigma=0.05$ are shown with blue solid curves and red dashed curves, respectively. Interestingly, the macroscopic magnetization of the noncanonical states near $\beta_{c}$ scales as
$M(t,\mu\sigma)=\mu^{1/8}M(t/\mu,\sigma).$
If we assume that $\sigma\sim 1/L$, where $L$ is the characteristic length of the system, such noncanonical states behave the same scaling with the finite-sized systems outside the thermodynamic limit (Fisher and Ferdinand, 1967; Ferdinand and Fisher, 1969; Binder, 1972; Privman and Fisher, 1984). The calculations related to FIG. 3 and the finite-size scaling can be found in Sec V in Supplementary Material (SM_, ).
Figure 3: Thermodynamic functions of canonical and noncanonical states of the two-dimensional Ising model. (a) Magnetization as a function of inverse temperature $\beta$ (b) Heat capacity as a function of inverse temperature $\beta$. The black dotted curves represent the canonical states. The blue solid curves and the red dashed curves represent the noncanonical states with $\Delta\beta/\beta=0.02$ and $\Delta\beta/\beta=0.05$, respectively.
Correlation from noncanonical states – The noncanonical distributions also reveal the information correlation between two events associated with $E_{m}$ and $E_{n}$. We consider a two-component system without interaction between the components, both of which have the same energy levels $E_{n}$. Unlike the canonical distribution, the joint distribution of the two components deviates from being a simple product of the distributions of each component, as discussed in Sec. VI in Supplementary Material (SM_, ). This is mathematically represented by the inequality $P_{mn}^{(\mathrm{nc})}(\beta)\neq P_{m}^{(\mathrm{nc})}(\beta)P_{n}^{(\mathrm{nc})}(\beta)$. Thus, the information entropy $S_{\mathrm{in}}(P_{n})=-\sum_{n}P_{n}\ln P_{n}$ for the subsystems exhibits a lack of additivity due to the noncanonical correlation, which is quantified by the mutual information $I\equiv S_{\mathrm{in}}(P_{m}^{(\mathrm{nc})})+S_{\mathrm{in}}(P_{n}^{(\mathrm{nc})})-S_{\mathrm{in}}(P_{mn}^{(\mathrm{nc})})$. We plot the mutual information of two correlated three-level systems as a function of the relative deviation of $\beta$ in FIG. 4.
@@
Figure 4: Mutual information between two-three level systems as a function of the relative deviation of inverse temperature.
@@
Actually, the noncanonicality of a black hole induces the information embedded in the correlation of its radiation, which has been considered a promising candidate for resolving the black hole information paradox. The question arises whether the noncanonical solutions to the EIE can describe such a noncanonical radiation field from a small Schwarzschild black hole. For a Schwarzschild black hole with temperature $\beta=8\pi M$, the deviation of $\beta$ is estimated as $\Delta\beta^{2}=\left|\beta^{2}/C\right|=8\pi$, where $C=-8\pi M^{2}$ is the heat capacity of the black hole (Hawking, 1975; Wald, 2001; Landau and Lifshitz, 2013; Carlip, 2014). From the noncanonical solutions of EIE, the emission rate (see Sec. VII in Supplementary Material (SM_, )) is given by
@@
which is valid in cases with small deviations. This result agrees with the previous findings from the microscopic theory for the radiation field near a small black hole (Parikh and Wilczek, 2000). The noncanonical state can account for the black hole information paradox (Zhang et al., 2009; Cai et al., 2016). It is emphasized that such noncanonical solutions presented in this letter arises from QTI. As black holes become smaller, the backaction of radiation grows stronger, leading to a tighter correlation between successive radiation events. This increased correlation causes the radiation state to deviate significantly from the traditional canonical thermal equilibrium state.
Conclusion – In conclusion, we have proposed an alternative approach to reconsider the foundations of statistical mechanics based on the QTI introduced in this Letter as a generalization of the path independence of work and heat. We established a quantum version of Carathéodory’s principle and derived the so-called EIEs, which can result in both canonical and noncanonical states. By considering detailed balance, we uniquely constructed the Boltzmann distribution from EIE for QTI as the foundation of statistical physics. In this sense, QTI not only reinforces the principles of statistical physics based on the experiment-driven principles of thermodynamics but also provides a new perspective on noncanonical thermodynamics grounded in information correlation. This work opens up new avenues for addressing long-standing problems in statistical thermodynamics, offering a more comprehensive framework for analyzing complex thermodynamic behaviors in small systems.
Acknowledgment – We are grateful to H. Dong, H. T. Quan, and Y. H. Ma for their helpful discussions. This work has been supported by the National Natural Science Foundation of China (NSFC) (Grant No. 12088101) and NSAF No.U2330401.